solve:
2x+3y=7
x+y=3

Answer by ewatrrr(24785)

 
Hi,
2x+3y=7
x+y=3
2x+3y=7
-2x-2y=-6 |multiplying thru by -2 t0 eliminate the x variable
y = 1 and x = 2 (x = -y+3 )
CHECKING our Answer***
2x+3y=7
4 + 3 = 7